### v1

# Sp2360: Sparse-view 360 Scene Reconstruction using Cascaded 2D Diffusion Priors

Abstract
We aim to tackle sparse-view reconstruction of a $360^{\circ}$ 3D scene using priors from latent diffusion models (LDM). The sparse-view setting is ill-posed and underconstrained, especially for scenes where the camera rotates $360$ degrees around a point, as no visual information is available beyond some frontal views focused on the central object(s) of interest. In this work, we show that pretrained 2D diffusion models can strongly improve the reconstruction of a scene with low-cost fine-tuning. Specifically, we present SparseSplat360 (Sp2360), a method that employs a cascade of in-painting and artifact removal models to fill in missing details and clean novel views. Due to superior training and rendering speeds, we use an explicit scene representation in the form of 3D Gaussians over NeRF-based implicit representations. We propose an iterative update strategy to fuse generated pseudo novel views with existing 3D Gaussians fitted to the initial sparse inputs. As a result, we obtain a multi-view consistent scene representation with details coherent with the observed inputs. Our evaluation on the challenging Mip-NeRF360 dataset shows that our proposed 2D to 3D distillation algorithm considerably improves the performance of a regularized version of 3DGS adapted to a sparse-view setting and outperforms existing sparse-view reconstruction methods in $360^{\circ}$ scene reconstruction. Qualitatively, our method generates entire $360^{\circ}$ scenes from as few as $9$ input views, with a high degree of foreground and background detail.
1 Introduction
Obtaining high-quality 3D reconstructions or novel views from a set of images has been a long-standing goal in computer vision and has received increased interest recently. Recent 3D reconstruction methods, such as those based on Neural Radiance Fields (NeRF) [38], Signed Distance Functions (SDFs) [64], or the explicit 3D Gaussian Splatting (3DGS) [30], are now able to produce photorealistic novel views of $360^{\circ}$ scenes. However, in doing so, they rely on hundreds of input images that densely capture the underlying scene. This requirement is both time-consuming and often an unrealistic assumption for complex, large-scale scenes. Ideally, one would like a 3D reconstruction pipeline to offer generalization to unobserved parts of the scene and be able to successfully reconstruct areas that are only observed a few times. In this work, we present a method to efficiently obtain high-quality 3D Gaussian representations from just a few views, moving towards this goal.
Standard 3DGS, much like NeRF, is crippled in a sparse observation setting. In the absence of sufficient observations and global geometric cues, 3DGS invariably overfits to training views. This leads to severe artifacts and background collapse already in nearby novel views due to the inherent depth ambiguities associated with inferring 3D structure from few-view 2D images. There exists a long line of work to improve performance of both NeRF and 3DGS in sparse novel view synthesis [14, 48, 32, 71, 78, 11, 62, 27, 42, 70]. While these works can reduce artifacts in sparsely observed regions, they are not able to fill in missing details due to the use of simple regularizations or weak priors.
In the setting of large-scale $360^{\circ}$ scenes, the problem is even more ill-posed and under-constrained. Much stronger priors are needed, such as those from large pre-trained 2D diffusion models [41, 45, 51, 49], capturing knowledge about typical structures in the 3D world. Recent approaches [35, 52, 69] add view-conditioning to these image generators by fine-tuning them on a large mixture of real-world and synthetic multi-view datasets. Leveraging these strong priors for optimization of radiance fields yields realistic reconstructions in unobserved areas of challenging $360^{\circ}$ scenes. In this work, we propose to forego augmenting a 2D diffusion model with additional channels for pose or context to make it 3D-aware. Instead, we perform low-cost fine-tuning of pretrained models to adapt them to specific sub-tasks of few-view reconstruction. This weakens the assumption of large-scale 3D training data, which is expensive to obtain.
We present SparseSplat360 (Sp2360), an efficient method that addresses the given task of sparse 3D reconstruction by iteratively adding synthesized views to the training set of the 3D representation. The generation of new training views is divided and conquered as the 2D sub-tasks of (1) in-painting missing areas and (2) artifact elimination. The in-painting model for the first stage is pre-trained on large 2D datasets and efficiently fine-tuned on the sparse views of the given scene. The artifact elimination network for the second stage is an image-to-image diffusion model, fine-tuned to specialize on removing typical artifacts appearing in sparse 3DGS. Thus, both stages utilize 2D diffusion models that are fine-tuned on small amounts of data. In each iteration, the models are conditioned on rendered novel views of the already existing scene, thus using 2D images as input and output, avoiding the requirement of training on large datasets of 3D scenes. In contrast to previous works, our method leverages stronger priors than simple regularizers and does not rely on million-scale multi-view data or huge compute resources to train a 3D-aware diffusion model.
In summary, our contributions are:
  • We present a novel systematic approach to perform sparse 3D reconstruction of $360^{\circ}$ scenes by autoregressively adding generated novel views to the training set.
  • We introduce a two-step approach for generating novel training views by performing in-painting and artifact removal with 2D diffusion models, which avoids fine-tuning on large-scale 3D data.
  • We introduce a sparse 3DGS baseline that improves reconstruction from sparse observation by applying regularization techniques without the need for pre-trained models.
  • We show that Sp2360 outperforms recent works in reconstructing large 3D scenes from as few as $9$ input views.
2 Related Work
2.1 Sparse-View Radiance Fields
Following the breakthroughs of NeRF [38] and 3D Gaussian Splatting [30] for inverse rendering of radiance fields, there have been many approaches to weaken the requirement of dense scene captures to sparse input views only. These methods can be categorized into regularization techniques and generalizable reconstruction priors.
Regularization Techniques
Fitting a 3D representation from sparse observations only is an ill-posed problem and very prone to local minima. In the case of radiance fields, this is typically visible as ’floaters’ during rendering of novel views. A classical technique for training with limited data is regularization. Many existing methods leverage depth from Structure-from-Motion [14, 48], monocular estimation [32, 71, 78, 11], or RGB-D sensors [62]. DietNeRF [27] proposes a semantic consistency loss based on CLIP [44] features. FreeNeRF [72] regularizes frequency range of NeRF inputs by increasing the frequencies of positional-encoding features in a coarse-to-fine manner. Moving closer to generative priors, RegNeRF [42] and DiffusioNeRF [70] maximize likelihoods of rendered patches under a trained normalizing flow or diffusion model, respectively.
Generalizable Reconstruction
In the case of very few or even a single view only, regularization techniques are usually not strong enough to account for the ambiguity in reconstruction. Therefore, another line of research focuses on training priors for novel view synthesis across many scenes. pixelNeRF [73] extracts pixel-aligned CNN features from input images at projected sample points during volume rendering as conditioning for a shared NeRF MLP. Similarly, many approaches [61, 7, 21, 34] define different NeRF conditionings on 2D or fused 3D features. Following the trend of leveraging explicit data structures for accelerating NeRFs, further priors have been learned on triplanes [26], voxel grids [18], and neural points [66]. Building upon the success of 3D Gaussian Splatting [30] and its broad applications to, e.g., surface [17, 25] or non-rigid reconstruction [36, 68, 12], recent methods like pixelSplat [6] and MVSplat [10] achieve state-of-the-art performance in stereo view interpolation while enabling real-time rendering. However, all of these works train regression models that infer blurry novel views in case of high uncertainty. Bridging the gap of generalizable and generative priors, GeNVS [5] and latentSplat [67] render view-conditioned feature fields followed by a 2D generative decoding to obtain a novel view.
2.2 Generative Priors
In case of ambiguous novel views, the expectation over all possible reconstructions might itself not be a reasonable prediction. Therefore, regression-based approaches fail. Generative methods, on the other hand, try to sample from this possibly multi-modal distribution.
Diffusion Models
In recent years, diffusion models [23, 15] emerged as the state-of-the-art for image synthesis. They are characterized by a pre-defined forward noising process that gradually destroys data by adding random (typically Gaussian) noise. The objective is to learn a reverse denoising process with a neural network that, after training, can sample from the data distribution given pure noise. Important improvements include refined sampling procedures [58, 29] and the more efficient application in a spatially compressed latent space compared to the high-resolution pixel space [49]. Their stable optimization, in contrast to GANs, enabled today’s text-to-image generators [41, 45, 51, 49] trained on billions of images [56].
2D Diffusion for 3D
While diffusion models have been applied directly on 3D representations like triplanes [57, 8], voxel grids [39], or (neural) point clouds [75, 37, 55], 3D data is scarce. Given the success of large-scale diffusion models for image synthesis, there is a great research interest in leveraging them as priors for 3D reconstruction and generation. DreamFusion [43] and follow-ups [63, 33, 9, 13, 59] employ score distillation sampling (SDS) to iteratively maximize the likelihood of radiance field renderings under a conditional 2D diffusion prior. For sparse-view reconstruction, existing approaches incorporate view-conditioning via epipolar feature transform [77], cross-attention to encoded relative poses [35, 52], or pixelNeRF [73] feature renderings [69]. However, this fine-tuning is expensive and requires large-scale multi-view data, which we circumvent with Sp2360.
3 Method
In this section, we describe our method in detail. The section begins with a general overview of Sp2360 in Sec. 3.1, outlining the autoregressive algorithm for training view generation. In the following sections, the individual parts of the system are introduced: the in-painting module in Sec. 3.2, the artifact removal procedure in Sec. 3.3, and the sparse 3DGS baseline method in Sec. 3.4.
3.1 Sp2360 for Sparse-View 3D Reconstruction
Given a sparse set of input images =$\mathcal{I}=\{\mathbf{I}_{1},\mathbf{I}_{2},...,\mathbf{I}_{M}\}$ with camera poses $\{\pi_{1},\pi_{2},...,\pi_{M}\}$, and a sparse point cloud $\mathbf{P}\in\mathbb{R}^{S\times 3}$, estimated by Structure-from-Motion (SfM) [53, 54], our goal is to obtain a 3D Gaussian representation of the scene, which enables the rendering of novel views from camera perspectives that are largely different from given views in $\mathcal{I}$. The tackled scenario of sparse-view inputs is extremely challenging, as the given optimization problem is heavily under-constrained and leads to severe artifacts if done naively.
Figure 1: Overview of Sp2360. We render 3D Gaussians fitted to our sparse set of $M$ views from a novel viewpoint. The image has missing regions and Gaussian artifacts, which are fixed by a combination of in-painting and denoising diffusion models. This then acts as pseudo ground truth to spawn and update 3D Gaussians and satisfy the new view constraints. This process is repeated for several novel views spanning the $360^{\circ}$ scene until the representation becomes multi-view consistent.
Algorithm 1 Sp2360 Algorithm
1:Sparse input image set $\mathcal{I}$, camera poses $\{\pi_{1},\pi_{2},...,\pi_{M}\}$, sparse point cloud $\mathbf{P}\in\mathbb{R}^{S\times 3}$
2:Set of 3D Gaussians $\mathcal{G}$
$\hat{\mathcal{I}}\leftarrow\mathcal{I}$
4:$\mathcal{G}\leftarrow$ Optimize Sparse 3DGS for $k$ iterations
5:for $N$ iterations do
6:     $\pi\leftarrow$ Sample novel camera pose.
7:     $\mathbf{I}\leftarrow R_{\pi}(\mathcal{G})$ Render from camera $\pi$
8:     $\mathbf{I}\leftarrow$ $(\mathbf{I})$
9:     $\mathbf{I}\leftarrow$ ArtifactRemoval$(\mathbf{I})$
10:     $\hat{\mathbf{I}}\leftarrow\ \hat{\mathbf{I}}\cup\{\mathbf{I}\}$
11:     $\mathcal{G}\leftarrow$ Optimize Sparse 3DGS for $k$ iterations
12:end for
An overview of Sp2360 is given in Fig. 1 and Alg. 1. The approach begins by optimizing a set of 3D Gaussians [30] to reconstruct the initial sparse set of input images $\hat{\mathcal{I}}=\mathcal{I}$. For this, we introduce a Sparse 3DGS baseline (c.f. Sec. 3.4), which combines best practices from previous works on NeRFs. The obtained representation serves as initial prior for $360^{\circ}$ reconstruction. Next, we autoregressively add new generated views to our training set: (1) we sample novel cameras and render novel views with artifacts and missing areas, make them look plausible by (2) performing in-painting (c.f. 3.2) and then (3) artifact removal (c.f. 3.3), before (4) adding them to set $\hat{\mathcal{I}}$ and continuing optimizing our 3D representation for $k$ iterations.
For training 3DGS in an iterative fashion, as outlined above, special precaution has to be taken to prevent overfitting on the initial views. An optimal schedule involves finding the number of iterations per cycle $k$ and the hyperparameters for the original 3DGS [30]. We provide a detailed evaluation in App. A.3. In the following, we will detail the in-painting and artifact removal stages of our pipeline.
3.2 In-Painting Novel Views
When using only a sparse observation set, many areas remain unobserved, leading to areas of no Gaussians in the 3D representation and zero opacity regions in some novel views (c.f. Fig.1). Regularization techniques cannot help with inferring details. Instead, we incorporate a generative in-painting diffusion model to fill such regions indicated by a binary mask $\phi$, which is obtained by rendering opacity from our current Gaussian representation $\mathcal{G}$ and binarizing it with a threshold $\tau$.
We fine-tune Stable Diffusion $2$ [49] to perform in-painting on our novel view renderings. The current training images $\hat{\mathcal{I}}$ are used as training data to adapt it to the current scene. For the fine-tuning technique, we get inspired by recent work [60] that uses LoRA [24] adapters for the UNet $\boldsymbol{\epsilon_{\theta}}$ and text encoder $c_{\theta}(y)$. Given images ^=$\hat{\mathcal{I}}=\{\mathbf{I}_{1},...,\mathbf{I}_{M}\}$, we create artificial masks $\{\phi_{1},...,\phi_{M}\}$ by creating random rectangular masks over each image and taking either their union or the complement of the union. Then, the adapter weights are fine-tuned using the following objective:
$\mathcal{L}=\mathbb{E}_{i\sim\mathcal{U}(M),y,\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{1}),t}\Big{[}\|\epsilon_{t}-\boldsymbol{\epsilon_{\theta}}(\mathbf{z}_{t};t,\phi_{i},c_{\theta}(y))\|_{2}^{2}\Big{]}\textnormal{,}$ (1)
where $\mathbf{z}_{t}$ is the diffused latent encoding of image $\mathbf{I}_{i}$ at step $t$. Here, $y$ is a simple text prompt - “A photo of [V]”, where [V] is a rare token like in DreamBooth [50] whose embedding is optimized for in-painting. For views $\mathbf{I}\notin\mathcal{I}_{i}$ (not from the original set), Eq 1 is only evaluated for regions that have a rendered opacity $>\tau$. This in-painting objective for fine-tuning enables our model to in-paint missing regions in a novel view rendering with details faithful to the observed $M$ views. At inference, $\boldsymbol{\epsilon_{\theta}}$ predicts the noise in $\mathbf{z}_{t}$ as:
$\hat{\epsilon_{t}}=\boldsymbol{\epsilon_{\theta}}(\mathbf{z}_{t};t,\phi_{i},c_{\theta}(y)),$ (2)
which is used to progressively obtain less noisy latents in $s$ DDIM [58] sampling steps starting from $t$. After passing the denoised latent $\hat{z}_{0}$ through a VAE decoder $\mathcal{D}$. we obtain the in-painted image $x_{0}$.
3.3 Removing Sparse-View Artifacts
The in-painting technique can fill in plausible details in low-opacity areas of novel view renders. However, it cannot deal with the dominance of blur, floaters, and color artifacts, which are not detected by the in-painting mask $\phi$. As such, we resort to a diffusion-based approach to learn how typical artifacts in 3D Gaussian representations from sparse views look like and how to remove them.
We fine-tune an image-conditioned diffusion model [4] to edit images based on short, user-friendly edit instructions. Here, the UNet $\boldsymbol{\epsilon_{\theta}}$ is trained to predict noise in $z_{t}$ conditioned on a given image $c_{I}$, in addition to the usual text conditioning $c_{\theta}(y)$. To enable this, additional input channels are added to the first convolutional layer of $\boldsymbol{\epsilon_{\theta}}$ so that $z_{t}$ and $\mathcal{E}(c_{I})$ can be concatenated. Weights of these channels are initialized to zero, whereas rest of the model is initialized from the pre-trained Stable Diffusion 1.5 [49]. For architectural design choices, we resort to those made by Instruct-Pix2Pix [4].
Dataset Creation
For training, we rely on a set $\mathcal{X}=\{(\mathbf{x}^{i},\mathbf{c}^{i},y^{i})_{i=1}^{N}\}$ of data triplets, each containing a clean image $\mathbf{x}_{i}$, an image with artifacts $\mathbf{c}_{i}$, and the corresponding edit instruction for fine-tuning $y_{i}$, to “teach” a denoiser to detect Gaussian artifacts and to generate a clean version of the conditioning image. For this, we build an artifact simulation engine comprising a full 3DGS model fitted to dense views, Sparse 3DGS fitted to few views, and camera interpolation and perturbation modules to use supervision of the full model at viewpoints beyond ground truth camera poses. The fine-tuning setup is illustrated in Fig. 2. For a given scene, we fit sparse models for $M\in\{3,6,9,18\}$ number of views. For larger $M$, we observe that there are very few artifacts beyond standard Gaussian blur. To have diversity in edit instructions, we start with a base instruction - “Denoise the noisy image and remove all floaters and Gaussian artifacts.” and ask GPT-4 [1] to generate 50 synonymous instructions. During training, each noisy, clean image pair is randomly combined with one of them.
Figure 2: Artifact removal fine-tuning. Pairs of clean images and images with artifacts are obtained from 3DGS fitted to sparse and dense observations, respectively, across $36$ scenes. These are combined with one of 51 synonymous prompts generated by GPT-4 from a base instruction. SD $1.5$ is then fine-tuned with a dataset of $10.5K$ samples for the Gaussian artifact removal task.
Training the Denoiser
Using our synthetically curated dataset $\mathcal{X}$, we fine-tune SD 1.5 as follows:
$\mathcal{L}=\mathbb{E}_{i\sim\mathcal{U}(N),\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I}),t}\Big{[}\|\epsilon_{t}-\boldsymbol{\epsilon_{\theta}}(\mathbf{z}^{i}_{t};t,\mathcal{E}(\mathbf{c}^{i}),c_{\theta}(y^{i}))\|_{2}^{2}\Big{]}$ (3)
where $\mathbf{z}_{t}^{i}$ is the encoded image $\mathbf{x}^{i}$ diffused with sampled noise $\epsilon$ at time step $t$. Thus, the model is fine-tuned to generate clean images $\mathbf{x}^{i}$, conditioned on artifact images $\mathbf{x}^{i}$ and text prompts $y^{i}$.
Generating Clean Renders
Given an in-painted rendering $\mathbf{x}_{0}$ and the base prompt $y=$"Denoise the noisy image and remove all floaters and Gaussian artifacts" at inference time, the fine-tuned artifact removal UNet $\boldsymbol{\epsilon_{\theta}}$ predicts the noise in latent $\mathbf{z}_{t}$ according to $t\sim\mathcal{U}[t_{min},t_{max}]$ as:
$\hat{\epsilon}_{t}=$ $\,\boldsymbol{\epsilon_{\theta}}(\mathbf{z}_{t};t,\varnothing,\varnothing)$ (4)
$+s_{I}\cdot(\boldsymbol{\epsilon_{\theta}}(\mathbf{z}_{t};t,\mathcal{E}(\mathbf{x}_{0}),\varnothing)-\boldsymbol{\epsilon_{\theta}}(\mathbf{z}_{t};t,\varnothing,\varnothing))$
$+s_{T}\cdot(\boldsymbol{\epsilon_{\theta}}(\mathbf{z}_{t};t,\varnothing,c_{\theta}(y))-\boldsymbol{\epsilon_{\theta}}(\mathbf{z}_{t};t,\mathcal{E}(\mathbf{x}_{0}),\varnothing))$
where $s_{I}$ and $s_{T}$ are the image and prompt guidance scales, dictating how strongly the final multistep reconstruction agrees with the in-painted render $\mathbf{x}_{0}$ and the edit prompt $y$, respectively. After $s$ DDIM [58] sampling steps, we obtain our final image by decoding the denoised latent.
3.4 Sparse 3DGS
The sparse 3DGS baseline serves as our starting point that already improves reconstruction quality over standard 3DGS. It is inspired by several recent works on sparse neural fields [14, 42] and 3DGS [71, 78, 32]. The baseline unifies depth regularization from monocular depth estimators and depth priors from pseudo views, while using specific hyperparameter settings, such as densification thresholds and opacity reset configurations for Gaussian splatting. It is outlined in detail in App. A.2 and evaluated individually in the ablation studies in Sec. 4.4.
4 Experiments
This section compares Sp2360 with state-of-the-art sparse-view reconstruction techniques. We also provide detailed ablation studies motivating specific design choices across different components of our approach. The setup for Iterative 3DGS (c.f. A.3) translates seamlessly to the iterative distillation procedure (c.f. A.4) with diffusion priors, and hence, we refrain from further analysis here.
4.1 Experimental Setup
Evaluation Dataset
We evaluate Sp2360 on the 9 scenes of the MipNeRF360 dataset [2], comprising 5 outdoor and 4 indoor scenes. Each scene has a central object or area of complex geometry with an equally intricate background. This makes it the most challenging $360^{\circ}$ dataset compared to CO3D [47], RealEstate 10K [76], DTU [28], etc. We retain the train/test split of the MipNeRF360 dataset, where every $8th$ image is kept aside for evaluation. To create $M$-view subsets, we sample from the train set of each scene using a geodesic distance-based heuristic to encourage maximum possible scene coverage (see supplement for details).
Fine-tuning dataset
We fine-tune the in-painting module only on the $M$ input views. For the artifact removal module, we train 3DGS on sparse and dense subsets of $360^{\circ}$ scenes from MipNeRF360 [2], Tanks and Temples [31] and Deep Blending [20] across a total of 36 scenes to obtain $\sim 10.5K$ data triplets. We train 9 separate artifact removal modules holding out the MipNeRF360 scene we want to reconstruct. On a single A100 GPU, fine-tuning the in-painting and artifact removal modules takes roughly $2$h and $1$h, respectively.
Baselines
We compare our approach against $7$ baselines. FreeNeRF [72], RegNeRF [42], DiffusioNeRF [70], and DNGaussian [32] are few-view regularization methods based on NeRFs or 3D Gaussians. ZeroNVS [52] is a recent generative approach for reconstructing a complete 3D scene from a single image. We use the ZeroNVS baseline introduced in ReconFusion [69], designed to adapt ZeroNVS to multi-view inputs. Conditioning the diffusion model on the input view closest to the sampled random view enables scene reconstruction for a general $M$-view setting. We also compare against 3DGS, the reconstruction pipeline for Sp2360, and against Sparse 3DGS, our self-created baseline. We cannot compare with ReconFusion [69] as neither their view splits nor code are publicly available, and their method builds on top of closed-source diffusion models. We expect to perform slightly worse due to weaker priors, but our approach is more lightweight and data-efficient.
Metrics
Due to the generative nature of our approach, we employ FID [22] and KID [3] to measure similarity of distribution of reconstructed novel views and ground truth images. We also compute two perceptual metrics - LPIPS [74] and DISTS [16] - to measure similarity in image structure and texture in the feature space. Despite their known drawbacks as evaluators of generative techniques [5, 52], we additionally provide PSNR and SSIM scores for completeness. Both favor pixel-aligned blurry estimates over high-frequency details, making them ill-suited to our setting.
4.2 Implementation Details
We implement our entire framework in Pytorch 1.12.1 and run all experiments on single $A100$ or $A40$ GPUs. We work with image resolutions in the 400-600 pixel range as this is closest to the output resolution of $512$ for both diffusion models. We set $\lambda_{1}=0.2$ (same as 3DGS) and vary $\lambda_{depth},\lambda_{pseudo}$ in $\{0.0,0.05,0.1\}$ for Sparse 3DGS. For in-painting, we set $\tau=0.8$ and fine-tune the in-painting module for 3000-5000 steps with LoRA modules of rank in $\{8,16,32\}$ (depending on $M$). We use a batch size of $16$ and learning rates of 2e-4 for the UNet and 4e-5 for the text encoder. The artifact removal module is trained for $2500$ iterations with batch size $16$ and learning rate 1e-4. To enable classifier-free guidance for both models, we randomly dropout conditioning inputs (text, mask, image, etc.) with probability of $0.1$ during training. The classifier-free guidance scales are set to $s_{I}=2.5$ and $s_{T}=7.0$. We use $t_{max}=0.99$ for both in-painting and artifact removal and linearly decrease $t_{min}$ for the in-painting module from $0.98$ to $0.90$, and from $0.98$ to $0.70$ for the artifact removal module. We sample both the in-painted and clean renders for $s=20$ DDIM sampling steps. We also linearly decay the weight of $\mathcal{L}_{sample}$ (c.f. A.4) from $1$ to $0.1$ over $30k$ iterations.
4.3 Comparison Results
We evaluate all baselines on our proposed splits for each scene. We report averaged quantitative results in Tables 1 and 2 and compare novel view rendering quality in Fig 3. We outperform all baselines for $9$-view reconstruction across all metrics and are second only to DiffusioNeRF on FID, KID, and DISTS for $3$ and $6$-view reconstruction. Unlike the baselines, our approach consistently improves with increasing $M$ across all metrics.
3DGS RegNeRF DNGaussian DiffusioNeRF ZeroNVS* Ours Ground Truth
Bicycle (9) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Treehill (9) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Garden (9) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Stump (9) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Kitchen (9) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Flowers (9) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Bonsai (6) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 3: Qualitative comparison of Sp2360 with few-view methods. Our approach consistently fairs better in recovering image structure from foggy geometry, where baselines typically struggle with “floaters” and color artifacts. We encourage the reader to refer to our supplemental 360° video, where the benefits of our method can be observed along a smooth trajectory.
Table 1: Quantitative comparison with state-of-the-art sparse-view reconstruction techniques on classical metrics. Despite being a generative solution, we outperform all baselines across all view splits on both pixel-aligned and perceptual metrics.
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Method 3-view 6-view 9-view 3-view 6-view 9-view 3-view 6-view 9-view
3DGS 10.288 11.628 12.658 0.102 0.141 0.190 0.709 0.661 0.605
FreeNeRF 9.948 9.599 10.641 0.124 0.129 0.145 0.682 0.679 0.668
RegNeRF 11.030 10.764 11.020 0.117 0.134 0.145 0.663 0.660 0.661
DNGaussian 9.867 10.671 11.275 0.124 0.137 0.179 0.754 0.730 0.729
DiffusioNeRF 10.749 11.728 11.430 0.093 0.116 0.112 0.709 0.678 0.654
ZeroNVS* 10.406 9.990 9.719 0.079 0.079 0.082 0.709 0.711 0.700
Ours 12.927 13.701 14.121 0.211 0.231 0.261 0.647 0.622 0.591
Table 2: Quantitative comparison with few-view reconstruction techniques on metrics suited for generative reconstruction. We are second only to DiffusioNeRF for $3$ and $6$-view reconstruction but achieve higher scores for $9$ views.
FID $\downarrow$ KID $\downarrow$ DISTS $\downarrow$
Method 3-view 6-view 9-view 3-view 6-view 9-view 3-view 6-view 9-view
3DGS 392.620 343.336 292.324 0.313 0.268 0.229 0.476 0.429 0.321
FreeNeRF 347.625 343.833 342.917 0.254 0.249 0.258 0.392 0.388 0.388
RegNeRF 362.856 347.045 349.043 0.291 0.266 0.247 0.399 0.403 0.404
DNGaussian 431.687 420.110 414.307 0.311 0.285 0.281 0.581 0.571 0.544
DiffusioNeRF 273.096 225.661 290.184 0.158 0.104 0.183 0.362 0.319 0.378
ZeroNVS* 351.090 335.155 337.457 0.283 0.282 0.290 0.437 0.429 0.428
Ours 318.470 283.504 230.565 0.273 0.229 0.162 0.384 0.357 0.315
4.4 Ablation Studies
In Tab LABEL:tab:ablation and Fig 4, we ablate the relative importance of each component towards $360^{\circ}$ reconstruction. We pick the garden scene and its $9$ view split for this study. We first show benefits of the regularization heuristics in Sparse 3DGS over native 3DGS. We attempt reconstruction using either the in-painting or artifact removal module and show that their combination works best for both novel view synthesis and final reconstruction. Interestingly, leaving out the artifact removal module results in the best PSNR and SSIM scores across all variants, emphasizing that blurry reconstructions are rewarded by classical metrics, whereas FID and KID favor sharp, realistic details in novel views. We remove the iterative update formulation and add in-painted clean renders for all novel views simultaneously. This variant performs worse than Sparse 3DGS, highlighting the need for autoregressive scene generation in our setting. Additionally, we try to supervise the renderings at novel views using the original 3DGS objective. However, this performs slightly worse than Eq 11. Note that an SDS-based formulation [52] is not applicable here due to the two-step nature of generative view synthesis.
Table 3: Ablation study on the $9$ view split of garden scene. Our combination of diffusion priors complements each other effectively. Without an iterative schedule to fuse novel views, our method fairs worse than a regularized baseline. Using the 3DGS objective for novel view renders leads to slightly worse performance on FID and KID.
Method FID $\downarrow$ KID $\downarrow$ DISTS $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$
3DGS 223.594 0.146 0.247 0.502 14.470 0.287
Sparse 3DGS 200.703 0.127 0.247 0.522 16.589 0.367
Sp2360 w/o Artifact R. 209.477 0.120 0.253 0.528 16.850 0.380
Sp2360 w/o In-painting 151.098 0.071 0.230 0.524 15.121 0.299
Sp2360 w/o Schedule 214.674 0.111 0.277 0.576 14.132 0.292
Sp2360 w/ $\mathcal{L}_{D-SSIM}$ 133.875 0.051 0.225 0.502 15.677 0.326
Sp2360 124.768 0.048 0.224 0.504 15.759 0.326
Sparse 3DGS w/o Artifact R. w/o In-painting w/o Schedule 3DGS Loss Ours Ground Truth
Render 1 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Render 2 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 4: Ablation Study on $9$-view reconstruction of garden scene. Our fine-tuned artifact removal module and iterative schedule contribute the most toward quality of the final reconstruction. 3D Gaussians from Sparse 3DGS act as suitable geometric prior in the absence of explicit view conditioning.
4.5 Scaling to More Views
In Fig 5, we analyze the relevance of diffusion priors with increasing $M$. We evaluate Sp2360 and 3DGS on view splits of increasing size with $M\in\{3,6,9,18,27,54,81\}$. For $M\leq 27$, our method consistently improves 3DGS’ generalization at novel views. We observe that as ambiguities resolve with increasing scene coverage, diffusion-based regularization becomes less important, and for $M\geq 54$, our method performs either equally or slightly worse across the 6 performance measures.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 5: Scalability of Sp2360 with input views. Our combination of fine-tuned diffusion priors improves performance of 3DGS up to $27$ input views of the bicycle scene, alleviating the need for dense captures.
4.6 Efficiency
Our approach focuses on limited use of multi-view data, minimal training times, and fast inference. We fit 3D Gaussians to our initial $M$ inputs inside $1$ hr and fine-tune the in-painting module in $\sim 2$ hrs. Sp2360 trains inside $30$ mins on a single A100 GPU, significantly faster than all baselines with NeRF backbones. FreeNeRF takes $\sim 1$ day while RegNeRF takes $>2$ days to train on a single A40 GPU. Both ZeroNVS* and DiffusioNeRF require $\sim 3$ hrs to distill 2D diffusion priors into Instant-NGP [40], whereas ReconFusion trains in 1 hour on 8 A100 GPUs  [^1] Source: ReconFusion authors. All 3 approaches use custom diffusion models that take several days to train on large-scale 3D datasets. On the contrary, we only need 10.5K samples to obtain a generalized artifact removal module that efficiently eliminates Gaussian artifacts and recovers image details in the foreground and background. We currently use a leave-one-out mechanism for the $9$ scenes in MipNeRF360 to train the artifact removal module for a particular MipNeRF360 scene. However, we only do this to have enough diverse data pairs across the $3$ datasets. This requirement can be easily alleviated using multi-view training data from other sources. Given the availability of a generalized artifact removal module, our entire pipeline can reconstruct a 360 scene in under $3.5$ hrs on a single A100 GPU. Courtesy of our 3D representation, we retain the real-time rendering capabilities of 3DGS.
5 Limitations & Conclusion
Limitations
Despite being a low-cost and efficient approach for reconstructing complex $360^{\circ}$ scenes from a few views, Sp2360 has certain limitations. Our approach is limited by the sparse geometry prior from an SfM point cloud, estimated from few views. For example, the point cloud for 9 views of the bicycle scene only has $628$ points. This prevents our method from achieving even higher fidelity and restricting artifacts in distant, ambiguous novel views. DUSt3R [65] is a recent stereo reconstruction pipeline that can potentially provide a stronger geometry prior.
Conclusion
We present Sp2360, a low-cost, data-efficient approach to reconstruct complex $360^{\circ}$ scenes from a few input images. We proposed a system that combines diffusion priors specialized for in-painting and Gaussian artifact removal to generate artificial novel views, which are iteratively added to our training image set. In our experiments we show that our approach approves over previous methods on the challenging MipNeRF360 dataset and illustrate the contributions of individual components in our ablation studies. For future work, we see the potential of our system to be improved with additional geometry cues from 3D vision foundation models.
Appendix A Appendix / Supplemental Material
In this section, we provide further details on certain heuristics and design choices of our pipeline not covered in the main paper. We begin with our strategy for creating $M$-view subsets of a scene (c.f A.1. We then provide details on the regularization introduced in our improved baseline to adapt 3DGS to sparse-view reconstruction (c.f A.2). We perform a toy experiment to determine the best choice of parameters for our iterative schedule (c.f. A.3). We end with some minor details on the distillation procedure to lift 2D diffusion priors to 3D Gaussians (c.f. A.4). Please also note our supplemental video, showing 360° reconstruction results.
A.1 Creating M-view Sparse Sets
3DGS relies on COLMAP for a sparse point cloud initialization, and COLMAP requires sufficient feature correspondence across the input images to produce a 3D consistent sparse scene geometry. To produce a subset where the views are far apart to have sufficient scene coverage and at the same time having enough feature correspondence among the images to produce a 3D consistent (albeit sparse) point cloud, we devise a well-defined strategy such that view subsets across different scenes follow a certain heuristic and are not just randomly picked.
We are given a set $\mathcal{X}=\{\mathbf{I}_{i},\pi_{i}\}_{i=1}^{N}$ of images and corresponding camera poses of a scene where $\mathbf{I}_{i}\in\mathbb{R}^{h\times w\times 3}$ and $\pi_{i}=[\mathrm{K}^{i}|\mathrm{R}^{i}|\mathrm{t}^{i}]\in SE(3)$. Here $K^{i}$ is the camera intrinsics matrix, $\mathrm{R}^{i}\in\mathbb{R}^{3\times 3}$ is the rotation matrix, and $t^{i}\in\mathbb{R}^{3\times 1}$ is the translation vector. From this dense set, we aim to create $M$-view sparse sets $(M<N)$$\{\mathbf{I}_{k},\mathrm{\pi}_{k}\}_{k=1}^{M}$ with a corresponding sparse point cloud $P_{M}\in\mathbb{R}^{S\times 3}$, obtained via COLMAP. To maximize the scene coverage of the $M$-view subset, we use the SE(3) (Special Euclidean group in three dimensions) representation to define a geodesic distance between a pair of camera viewpoints. SE(3) represents all rigid transformations (translations and rotations) that can be applied to a three-dimensional space while preserving distances and angles. Mathematically, SE(3) is defined as:
$SE(3)=\{(\mathrm{R},t)|\mathrm{R}\in SO(3),t\in\mathbb{R}^{3}\}$ (5)
Here, $SO(3)$ is the special orthogonal group in three dimensions, representing all possible 3D rotations without reflection or scaling. To quantify “closeness” between 2 SE(3) elements $\mathrm{\pi}_{1}$ and $\mathrm{\pi}_{2}$, we use the geodesic distance that measures the shortest path between 2 viewpoints on a curved surface - in this case, the minimum distance along a curve in SE(3) that connects $\mathrm{\pi}_{1}$ and $\mathrm{\pi}_{2}$. One common approach of measuring the geodesic distance involves separating the rotational and translational components and combining individual distances through a weighted average. If $\mathrm{R_{1}}$, $\mathrm{T_{1}}$ and $\mathrm{R_{2}}$, $\mathrm{T_{2}}$ are the rotation and translation matrices corresponding to $\mathrm{\pi}_{1}$ and $\mathrm{\pi}_{2}$ respectively, the geodesic distance can be defined as:
$\mathcal{D}_{geodesic}=\lVert\textnormal{Rodrigues}(\mathrm{R}_{rel})\rVert+w_{T}\lVert\mathrm{T_{1}}-\mathrm{T_{2}}\rVert$ (6)
where the Rodrigues formula determines the relative angle of rotation $\theta$ between 2 matrices $\mathrm{R_{1}}$ and $\mathrm{R_{2}}$ by measuring the angle of the difference rotation matrix $\mathrm{R}_{rel}=\mathrm{R_{1}}\mathrm{R_{2}}^{T}$ as follows:
$\textnormal{Rodrigues}(\mathrm{R}_{rel})=\mathcal{\theta}=\cos^{-1}\left(\frac{\text{trace}(R_{1}R_{2}^{T})-1}{2}\right)$ (7)
$w_{T}$ is a weighting factor. We set $w_{T}=0.1$ for all our experiments as we usually find that rotational distance has more impact on scene coverage given an $M$-view constraint.
(a) $M=9$ in the $treehill$ scene
(b) $M=18$ in the $stump$ scene
Figure 6: Plot of maximum geodesic distance across training viewpoints and SfM point cloud size vs. $n$ the index of the $n^{th}$ closest viewpoint from the available training stack. The locations of the maxima in the blue plots are used to create our view subset of size $M$.
Figure 7: We find that the maximum geodesic distance between all possible pairs of cameras in a view subset directly correlates with how much of a $360^{\circ}$ scene we can observe. Here we show two farthest views (in terms of $\mathcal{D}_{geodesic}$) of the treehill scene for $M=9$.
[…]
An example plot is shown in Fig 6 where we observe the change in $\mathcal{D}^{\mathrm{\pi}}(n)$ and the point cloud size $P_{M}$ as we steadily increase $n$ by 1. For $M=9$ in the treehill scene (left), we pick the subset corresponding to $n=17$, the global maxima for $\mathcal{D}^{\mathrm{\pi}}(n)$. Similarly, for $M=18$ in the stump scene (right), we pick the subset corresponding to $n=31$. Visually, in Fig 7, we observe that this heuristic gives us considerable scene coverage.
A.2 Sparse 3DGS
Inspired by recent works [71, 78], we introduce 4 different regularization techniques to prevent 3DGS from overfitting to the observed few views.
Tunable Hyperparameters
Similar to [11], we observe that the periodic resetting of Gaussian opacities, designed to restrict over-reconstruction in a dense setting, harms generalization to novel views in a sparse setting. Either several Gaussians are removed through pruning or get trapped in local minima of the optimization process, not receiving enough updates to match the given scene. Visually, we observe much lower artifacts in novel views for extremely sparse settings ($M\leq 18$) if opacities $\alpha$ are optimized with no resets during training. The densification threshold for positional gradients - $\tau_{pos}$ is another hyperparameter that controls the overall number of Gaussians. Experimentally, we observe varying $\tau_{pos}$ between 2-5 times the default value for dense setting in 3DGS, leads to smoother geometry and prevents overfitting in a sparse setting for $M\leq 54$. We show visual evidence in Fig 8 and 9 below.
(a) No opacity ($\alpha$) reset for $30k$ iterations.
(b) opacity ($\alpha$) reset every $3k$ iterations.
Figure 8: Periodic reset of the opacity ($\alpha$) of Gaussians has a detrimental effect on novel view synthesis in a sparse reconstruction setting. For example, scene details are more clearly observed in the left image, where there is no $\alpha$ reset for $30k$ iterations. ($M=9$)
(a) $\tau_{pos}=0.0002$
(b) $\tau_{pos}=0.001$
Figure 9: Effect of densification threshold ($\tau_{pos}$) for $M=9$ in the $bicycle$ scene. A higher value of $\tau_{pos}$ in a view-constrained scenario limits artifacts and favours smoother geometry.
Depth Regularization
Inspired by depth regularization techniques in NeRF [14, 42] and 3DGS literature [32], we introduce dense depth priors from pretrained monocular depth estimators (MDEs) [46]. Due to the severe sparsity of the initial point cloud $P_{M}$, we observe that this regularization is essential to make Gaussians grow towards a reasonable approximation of the true scene geometry. If $\mathcal{D}_{est}$ is the estimated monocular depth from a training view, and $\mathcal{D}_{ras}$ is the rendered depth, we employ the Pearson correlation coefficient (PCC) as a loss function to align the two depth map distributions [71, 78]. Specifically, we add a third loss term to the 3DGS optimization as:
$\mathcal{L}_{depth}=1-PCC(\hat{\mathcal{D}}_{ras},\hat{\mathcal{D}}_{est})=1-\frac{Cov(\hat{\mathcal{D}}_{ras},\hat{\mathcal{D}}_{est})}{\sqrt{Var(\hat{\mathcal{D}}_{ras})Var(\hat{\mathcal{D}}_{est})}}$ (8)
where $\hat{\mathcal{D}}_{ras}$, $\hat{\mathcal{D}}_{est}$ are obtained after normalizing $\mathcal{D}_{ras}$ and $\mathcal{D}_{est}$ to the range $[0,1]$.
Pseudo Views
As a final regularization, we generate pseudo views using camera interpolation and employ depth regularization as above to improve generalization in unobserved areas. We synthesize novel views on a path by interpolating between a viewpoint $\pi_{1}$ and its closest view $\pi_{2}$ in SE3 space. Specifically, we sample $\mathbf{t}$ on a B-spline fitted between the translation matrices and estimate an averaged orientation $\mathbf{R}$ through slerp interpolation on the rotation matrices. This yields an augmented viewpoint $\hat{\pi}$ with shared camera intrinsics and camera extrinsics $(\mathbf{R}|\mathbf{t})$. Without a ground truth image for $\hat{\pi}$, we only enable depth regularization on the rendered RGB image. Overall, the optimization objective becomes:
$\mathcal{L}=(1-\lambda_{1})\mathcal{L}_{1}+\lambda_{1}\mathcal{L}_{D-SSIM}+\lambda_{depth}\mathcal{L}_{depth}+\lambda_{pseudo}\mathcal{L}_{pseudo}$ (9)
We switch on $\mathcal{L}_{pseudo}$ after 2000 iterations when the optimization has stabilized, and Gaussians have learned at least a coarse scene geometry. We call this improved baseline Sparse 3DGS. Visually, we observe in Fig 10 that depth regularization through both ground truth and synthesized views helps 3DGS learn plausible geometry of the scene - like the structure of the bicycle and bench while smoothening out a lot of the grainy noise in the background.
(a) $\lambda_{depth}=0.05,\lambda_{pseudo}=0.05$
(b) $\lambda_{depth}=\lambda_{pseudo}=0.0$
Figure 10: Effect of $\mathcal{L}_{depth}$ and $\mathcal{L}_{pseudo}$ for $M=9$ in $bicycle$ scene. Enabling a depth-based regularization loss during training helps resolve depth ambiguities and learn true scene geometry at nearby novel viewpoints.
A.3 Iterative 3DGS
Our iterative schedule fuses clean, in-painted novel view renders with the 3D Gaussians fitted by Sparse 3DGS. This part of our method is inspired, in part, by Instruct NeRF2NeRF [19]. However, unlike their setting, we do not have ground truths for novel views to condition our diffusion models, making it more challenging. We propose to find the best combination of parameters for our algorithm in a simplified setting.
We first fit 3D Gaussians to the $9$ view split of the bicycle scene and then autoregressively sample closest novel views to fuse corresponding ground truth image with existing scene Gaussians. We argue that this variant should closely match the performance of a full model fitted to dense views. Both models see the same training views but with different initializations - SfM point cloud of the entire scene for the full model vs 3D Gaussians fitted to $M$ views for Iterative 3DGS. The full model sees all training views simultaneously, while the iterative version does so incrementally, adjusting the number and position of Gaussians with information from each novel view.
We fuse the heldout $N-M$ novel views over 30k iterations, sampling $m$ views at a time where $m\in\{1,2,3\}$. After each incremental update of the training stack, the Gaussians are optimized for $n_{k}$ iterations where $\{n_{k}\}_{k=1}^{\lceil\frac{N-M}{m}\rceil}$ is sampled from a schedule, which is a function of the total iterations (30k) and the number of novel views $N-M$. Specifically, $n_{k}$’s satisfy the following constraint:
$\sum_{k=1}^{\lceil\frac{N-M}{m}\rceil}n_{k}$ $=30000$ (10)
$n_{k+1}$ $=f(k)\cdot n_{k}\hskip 20.00003ptk\in\{1,2,\ldots,\lceil\frac{N-M}{m}\rceil-1\}$
In the simplest case, when we have a constant number of iterations per train stack update, $f(k)=1$ and $n_{k}=\left\lfloor\frac{30000}{\frac{N-M}{m}}\right\rfloor$. For linear and quadratic schedules, $f(k)=a\cdot k$ or $f(k)=a^{2}\cdot k$, where $a$ is calculated from the summation constraint (Eq 10). A generic solver for Eq 10 typically finds the largest $a$ for which $\sum_{k}n_{k}\leq 30000$ is satisfied, and the residual iterations are added to one of the $n_{k}$’s, usually $n_{\lceil\frac{N-M}{m}\rceil}$.
Table 4: Ablation Study on design choices of Iterative 3DGS for $M=9$ in $bicycle$ scene. We fuse ground truth images for heldout novel views and attempt to match the performance of a full model.
[…]
$m=1$ + constant 21.13 0.51 0.39
$m=2$ + constant 21.87 0.52 0.36
$m=3$ + constant 21.29 0.53 0.39
$m=2$ + cosine 21.01 0.52 0.37
$m=2$ + linear 22.26 0.57 0.34
$m=2$ + quadratic 22.43 0.58 0.33
$m=2$ + quadratic + no $\alpha$-reset 20.37 0.48 0.39
$m=2$ + quadratic + $\eta=0.95$ 22.72 0.60 0.32
$m=2$ + quadratic + $\eta=0.97$ ($\mathcal{O}$) 22.95 0.63 0.29
Full Model 24.34 0.73 0.22
Findings
We summarize the findings of our toy experiment in Tab 4. We experiment with constant, linear, quadratic, and cosine schedules and find that a quadratic schedule gets us closest to the performance of the full model. We find $m=2$ gives better performance compared to adding just 1 novel viewpoint at a time. We hypothesize that adding 2 viewpoints at a time provides an implicit depth prior that better guides incremental optimization of the Gaussians. Adding $m>2$ viewpoints does not provide any further improvement. Each Gaussian has a 3D scaling vector $s$ as one of its attributes that determines the scaling of the ellipsoid in 3 axial directions in world space. Our experiments show that shrinking $s$ by a scale adjustment factor $\eta<1$, once every set of viewpoint(s) is added, impacts generalization to novel views. Intuitively, some of the Gaussians, fitted to sparse views, would have grown towards large volumes to create over-smoothed geometry in empty space. So, shrinking the extent of the ellipsoids, as further observations are introduced, gives more incentive to the adaptive control module to clone Gaussians in under-reconstructed regions and reconstruct small-scale geometry better.
Lastly, we find that regularizations introduced in Sparse 3DGS are no longer needed during the iterative process due to the growing set of observations and diminishing ambiguities. As such, setting $\tau_{pos}=0.0002$, resetting opacity($\alpha$) every $3k$ iterations, and removing $\mathcal{L}_{depth}$, $\mathcal{L}_{pseudo}$ gets us closest to the performance of the full model. We call the best-performing variant our oracle $\mathcal{O}$, which gives us the experimental maximum for performance on the held-out test set. $\mathcal{O}$ also has similar storage requirements as the full model ($3.82M$ v $3.80M$ Gaussians), meaning the iterative 3DGS does not lead to a drop in FPS at inference time.
A.4 Distilling 2D priors to 3D
Our diffusion priors infer plausible detail in unobserved regions and the iterative update algorithm incrementally grows and updates scene Gaussians by fusing information at novel viewpoints. We initialize the scene with 3D Gaussians fitted by Sparse 3DGS to $M$ input views, autoregressively sample closest novel viewpoints (based on SE3 distance), obtain pseudo ground truths for novel views using our combination of diffusion priors, add them to the training stack and optimize for certain iterations (determined by the schedule in Sec A.3). At every iteration, we sample either an observed or unobserved viewpoint from the current training stack. We optimize Gaussian attributes using the 3DGS objective for known poses and the SparseFusion [77] objective for novel views:
$\mathcal{L}_{sample}(\psi)$ $=\mathbb{E}_{\pi,t}\Big{[}w(t)(\|I_{\pi}-\hat{I}_{\pi}\|_{1}+\mathcal{L}_{p}(I_{\pi},\hat{I}_{\pi}))\Big{]}$ (11)
where $\mathcal{L}_{p}$ is the perceptual loss [74], $w(t)$ a noise-dependent weighting function, $I_{\pi}$ is the 3DGS render at novel viewpoint $\pi$, and $\hat{I}_{\pi}$ is the inpainted, clean render version of $I_{\pi}$ obtained with our cascaded diffusion priors.